OpenStudy (astrophysics):

How do I set up this integral, under the surface z= 1+x^2y^2 and above the region enclosed by x = y^2 and x = 4.

OpenStudy (astrophysics):

\[z=1+x^2y^2\] basically this part :P

OpenStudy (astrophysics):

Or is their a way to do this, without having to draw it out?

OpenStudy (astrophysics):

@ganeshie8

ganeshie8 (ganeshie8):

you want to setup the volume integral is it ?

OpenStudy (astrophysics):

Yup

ganeshie8 (ganeshie8):

try this \[\int\limits_{-2}^2~~\int\limits_{y^2}^4~\int\limits_{0}^{1+x^2y^2}~1 ~dzdxdy\]

OpenStudy (astrophysics):

How did you get the -2 to 2 ?

ganeshie8 (ganeshie8):

solve x=y^2 and x=4 before even doing that, sketch the curves in xy plane maybe..

OpenStudy (astrophysics):

The integral IS \[\int\limits_{-2}^{2}\int\limits_{y^2}^{4} (1+x^2y^2)dxdy\] but I don't know how to get the -2 to 2 limits

ganeshie8 (ganeshie8):

sketch below curves in xy plane x = y^2 x = 4

OpenStudy (astrophysics):

That's it haha?

ganeshie8 (ganeshie8):

after sketching if it becomes obvious, then yes thats it :)

OpenStudy (astrophysics):

Oh I see, my problem was I was sketching y=x^2 xDDDD

OpenStudy (astrophysics):

Thanks!! :D

ganeshie8 (ganeshie8):

if you want it more challenging, you may try setting up `dydx` instead of `dxdy` :P

OpenStudy (astrophysics):

Loool I'll try that for fun